Think of a number multiply it by 3 then add 3 I get 12 what is the number

Answer by staceyd88(1)

The answer is 3.

let X represent the number your looking for, and write the equation out accordingly.
3(x)+3=12
3x+3=12
3x+3-3=12-3 subtract 3 on both sides
3x=9 now divide both sides by 3 to get x alone
3x/3=9/3 x=3
Now plug it back into the equation to see if its the answer your looking for.
(3)(3)+3=12
9+3=12
12=12